Bob's lawn-mowing service is a profit-maximizing, competitive firm. Bob mows lawns for $27 each. His total cost each day is $280, of which $30 is a fixed cost. He mows 10 lawns a day.

In the short run, Bob should( ). In the long run, Bob should( )the industry.
Answer #1

Answer

In a short run a fixed cost is not recoverable and thus person produces if in the short run it will recover his total variable cost i.e. Total Revenue > Total Variable cost

Here Total Revenue = Price*Quantity = 27*10 = 270.

Total Variable Cost = Total Cost - Total fixed cost = 280 - 30 = 250

Thus, Total Revenue > Total Variable cost. Hence, Bob Should produce in the Short Run.

In the long run, a fixed cost is recoverable and thus If a firm incurs losses then firm will exit the market.

Here, Profit = Total Revenue - Total Cost = 270 - 280 = -10 < 0.

Hence, Bob should exit the market in the long run.

Hence,In the short run, Bob should Produce and In the Long run Bob Should leave the industry.